# Historically Relevant Event Structuring for Temporal Knowledge Graph Reasoning

### v1

@@
Abstract.
Temporal Knowledge Graph (TKG) reasoning focuses on predicting events through historical information within snapshots distributed on a timeline. Existing studies mainly concentrate on two perspectives of leveraging the history of TKGs, including capturing evolution of each recent snapshot or correlations among global historical facts. Despite the achieved significant accomplishments, these models still fall short of (1) investigating the influences of multi-granularity interactions across recent snapshots and (2) harnessing the expressive semantics of significant links accorded with queries throughout the entire history, especially events exerting a profound impact on the future. These inadequacies restrict representation ability to reflect historical dependencies and future trends thoroughly. To overcome these drawbacks, we propose an innovative TKG reasoning approach towards Historically Relevant Events Structuring ($\mathsf{HisRES}$). Concretely, $\mathsf{HisRES}$ comprises two distinctive modules excelling in structuring historically relevant events within TKGs, including a multi-granularity evolutionary encoder that captures structural and temporal dependencies of the most recent snapshots, and a global relevance encoder that concentrates on crucial correlations among events relevant to queries from the entire history. Furthermore, $\mathsf{HisRES}$ incorporates a self-gating mechanism for adaptively merging multi-granularity recent and historically relevant structuring representations. Extensive experiments on four event-based benchmarks demonstrate the state-of-the-art performance of $\mathsf{HisRES}$ and indicate the superiority and effectiveness of structuring historical relevance for TKG reasoning.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Temporal reasoning
1. Introduction
@@
Temporal Knowledge Graphs (TKGs) (Cai et al., 2022) have revolutionized the limitations of static KGs by considering temporal information. Specifically, TKGs advance beyond traditional triple facts, enhancing them into (subject, relation, object, timestamp) quadruples, thereby bridging the divide between static KGs and dynamic characteristics. This enrichment facilitates capturing the dynamic evolution of real-world events and holds significant practical relevance. However, TKGs still suffer from incompleteness and encounter unique issues due to their distinct temporal nature. This pressing task in TKGs can be divided into interpolation and extrapolation, involving reasoning missing facts (events) at past and future timestamps according to historical knowledge, respectively. Notably, this paper primarily focuses on the extrapolation task that has been widely explored in recent years. It expands the downstream application domains of KGs to include event prediction (Zhou et al., 2023), social network analysis (Cai et al., 2018), sensor fault detection (Cai and Zheng, 2020), etc.
As outlined in prior research (Han et al., 2021; Li et al., 2021b), TKGs manifest in two primary forms. One form is a multi-relational graph akin to static KGs but enhanced with timestamped relation edges to convey dynamic characteristics, while the other can be represented as sequences of snapshots or subgraphs partitioned by timeline. In the latter form, each snapshot encapsulates concurrent events transpiring at specific timestamps, offering a more intuitive depiction of the evolutionary patterns of events.
In the extrapolation task, several works introduce novel perspectives that emphasize different aspects of historical information within TKGs. According to the views and ideas, they can be divided into two categories: (a) tallying statistics of historical events and leveraging repetitive patterns to predict future facts (Zhu et al., 2021; Xu et al., 2023); (b) capturing concurrent and evolutionary correlations from the most recent snapshots (Jin et al., 2020; Li et al., 2021b, 2022a; Liu et al., 2023; Liang et al., 2023; Zhang et al., 2023a). In detail, the (a) category utilizes simple linear transformations as encoders and generates sparse matrices from repetitive statistics to mask predictions, while the (b) category adopts Graph Neural Networks (GNNs) to model the structural dependencies among concurrent facts, and recurrent units such as Gated Recurrent Unit (GRU) to capture the sequential evolution of snapshots.
Based on the (b) category, few innovative models (Li et al., 2022c; Zhang et al., 2023b; Chen et al., 2024) additionally incorporate information from the entire history to achieve more comprehensive learning of TKGs. Within them, TiRGN (Li et al., 2022c) incorporates a global historical encoder like CyGNet (Zhu et al., 2021) to constrain the prediction scope, HGLS (Zhang et al., 2023b) establishes connections between common entities across different timestamps to capture multi-hop correlations among globally historical events, LogCL (Chen et al., 2024) constructs a graph consisting of historical facts relevant to queries to obtain rich semantic information. However, despite these advancements, the above studies still fall short of comprehensively exploring historically relevant events in TKGs. It can be analyzed into two specific challenges: (1) investigating the influences of multi-granularity interactions across recent snapshots and (2) harnessing the expressive semantics of significant links accorded with queries throughout the entire history, especially events exerting a profound impact on the future.
Figure 1. An illustration of extrapolation over TKGs.
Specifically, the first noteworthy challenge arises due to the lack of explicit and flexible judgment on multi-granularity correlations between facts or entities across adjacent snapshots. Since many sequentially occurring real-world events at adjacent timestamps are causally linked, the above-mentioned deficiency leads to the loss of crucial dependencies between such events from adjacent timestamps. It could originally provide valuable clues for TKG reasoning. While RPC (Liang et al., 2023) and LogCL (Chen et al., 2024) involve inter-snapshot information and assign different weights to distinguish contributions of recent snapshots, they remain limited to the snapshot level and may overlook specific links within snapshots. A vivid example from ICEWS (Boschee et al., 2015) dataset is illustrated in Figure 1, (North_America, Host_a_visit, Business_(Africa)) is caused by (Barack_Obama, Consult, North_America) at the previous time. Instead of complex evolutionary patterns, we can capture such correlations with simple 2-hop links (denoted in red) through structuring and learning events inter-snapshot.
For the second significant challenge, prior approaches tend to neglect valuable information intrinsic in the structure of globally relevant events, such as long-term semantics and periodic interactions, leading to inadequate representations. In Figure 1, it is evident that the triple ($e_{o}$, $r_{0}$, $e_{1}$) in $\mathcal{G}_{0}$ holds profound implications for future events. By identifying significant facts relevant to queries like this and structuring links (denoted in blue) throughout the entire history, we can effectively leverage crucial events from a global perspective. Although HGLS (Zhang et al., 2023b) proposes to connect common entity in each snapshot through additional edges to capture entity dependencies, it may incorporate redundant information from distant timestamps while necessitating more hops and layers for modeling neighboring interactions. LogCL (Chen et al., 2024) provides a similar idea for integrating historical facts relevant to queries, but it lacks the means to prioritize key historical links, leading to reduced efficiency.
@@
To tackle these aforementioned challenges, this paper proposes a novel method that specializes in TKG reasoning with Historically Relevanct Event Structuring ($\mathsf{HisRES}$). Specifically, $\mathsf{HisRES}$ adopts an encoder-decoder architecture, focusing on two critical aspects of TKG reasoning: capturing the structural and temporal patterns within recent snapshots and understanding the impact of relevant events throughout the entire history. In terms of recent history, $\mathsf{HisRES}$ employs a multi-granularity evolutionary encoder (§3.2), which models concurrent interactions and evolution within each snapshot and further integrates adjacent snapshots to learn temporal dependencies using multi-hop links (addressing the first challenge). To explicitly capture structural correlations across the entire history (addressing the second challenge), $\mathsf{HisRES}$ introduces a global relevance encoder, structuring historical facts directly relevant to the query set. Specially, we propose an innovative convolution-based graph attention network (ConvGAT) to efficiently emphasize important links of historical relevance for future prediction. Furthermore, $\mathsf{HisRES}$ incorporates a self-gating mechanism (§3.3) to adaptively fuse entity representations from different granularities or encoders, achieving a comprehensive understanding by considering various historical information (i.e., global and local, intra-snapshot and inter-snapshot) of TKGs, leading to accurate predictions. Extensive experiments on four benchmarks demonstrate the superiority and effectiveness of $\mathsf{HisRES}$ in TKG reasoning, particularly in structuring historical relevance.
@@
  • We propose a novel TKG reasoning method termed $\mathsf{HisRES}$, which sufficiently discovers structural and temporal dependencies within TKGs through historically relevant event structuring.
@@
  • Through the adaptive fusion of multi-granularity evolutionary and global relevance encoding, $\mathsf{HisRES}$ can comprehensively integrate representations of TKGs throughout historical correlations.
@@
  • Extensive experiments on four event-based TKG datasets demonstrate that $\mathsf{HisRES}$ outperforms state-of-the-art TKG reasoning baselines, showcasing its effectiveness and superiority.
2. Related Works
@@
KG reasoning (Tian et al., 2021) has been extensively studied in recent years, making significant contributions to downstream applications. Traditional reasoning deals mainly with static KGs such as the widely used FB15k-237 (Toutanova et al., 2015) and WN18RR (Dettmers et al., 2018). In contrast, TKG reasoning is tailored for time-sensitive data, adept at capturing evolutionary patterns, and discovering future facts within event-based datasets characterized by temporal features.
Figure 2. An illustration of $\mathsf{HisRES}$ model architecture.
2.1. Static Knowledge Graph Reasoning
@@
Early research on static KGs focuses on modeling entities and relations in various representation spaces. TransE (Bordes et al., 2013) introduces a simple translation assumption $s+r\approx o$, while DistMult (Yang et al., 2015) learns relational semantics through embeddings from a bilinear objective. In the complex vector space, ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019) employ Hermitian dot products and relation rotations, respectively. With the advent of convolutional neural networks, approaches such as ConvE (Dettmers et al., 2018) and ConvTransE (Shang et al., 2019) utilize convolutional operations to fuse entity-relation embeddings. Recent advancements aim at learning more expressive structures with Graph Convolutional Networks (GCN) (Kipf and Welling, 2017). R-GCN (Schlichtkrull et al., 2018) extending GCNs with relation-specific block-diagonal matrices. CompGCN (Vashishth et al., 2020) integrates multi-relational GCN methods with various composition operations, while KBGAT (Nathani et al., 2019) introduces an attention-based mechanism to capture neighboring semantics. However, these static approaches fail to incorporate temporal information effectively, making it hard to handle dynamic facts or learn evolutionary patterns.
2.2. Temporal Knowledge Graph Reasoning
@@
According to the occurrence moment of predicted events (i.e., known or future timestamps), TKG reasoning can be roughly divided into two types: interpolation and extrapolation. Notably, this paper focuses on the extrapolation task.
@@
Assuming we know all facts before $t_{T}$ (including $t_{T}$), the extrapolation task aims to predict future facts at timestamps $t>t_{T}$ based on historical TKG sequences. Various approaches have been explored in recent years to address it. xERTE (Han et al., 2021) designs a subgraph sampling module based on temporal information and incorporates time within relation embeddings. TITer (Sun et al., 2021) and CluSTER (Li et al., 2021a) employ reinforcement learning to generate query-specific paths composed of historical entities. Besides, numerous studies can be categorized into two main groups: those that focus on repeating statistical historical facts, and those that model the interactions among historical events. In the first category, CyGNet (Zhu et al., 2021) utilizes multi-hot vector as a historical vocabulary that records facts from the past and masks prediction. Building upon this, CENET (Xu et al., 2023) introduces contrastive learning and learnable classifiers to differentiate historical events. In the latter category, RE-NET (Jin et al., 2020) and RE-GCN (Li et al., 2021b) employ GCNs to model concurrent structures and utilize recurrent units to learn entity or relation evolution patterns. On this basis, CEN (Li et al., 2022a) considers historical length-diversity and time-variability. TiRGN (Li et al., 2022c) adopts a global history encoder like CyGNet to identify entities related to historical or unseen facts and a time-guided decoder. RETIA (Liu et al., 2023) and RPC (Liang et al., 2023) introduce the concept of line graphs to learn the interaction of relations. L2TKG (Zhang et al., 2023a) addresses the missing associations and creates links between unconnected entities to obtain more comprehensive representations. Some novel insights attempt to capture global structural correlations through additional graphs. HGLS (Zhang et al., 2023b) explicitly captures long-term dependencies by connecting the same entity at each timestamp, while LogCL (Chen et al., 2024) proposes an entity-aware attention mechanism to flexibly adopt facts relevant to queries and utilizes contrastive learning to improve robustness. Nevertheless, none of these methods consider the multi-granularity interactions among the most recent snapshots, which are crucial for predicting future facts. Nor do they efficiently learn correlations from the entire history, resulting in somewhat excessive consumption, such as increasing the scale of the graph or introducing unnecessary facts.
3. Method

### v2

@@
Abstract
Temporal Knowledge Graph (TKG) reasoning focuses on predicting events through historical information within snapshots distributed on a timeline. Existing studies mainly concentrate on two perspectives of leveraging the history of TKGs, including capturing evolution of each recent snapshot or correlations among global historical facts. Despite the achieved significant accomplishments, these models still fall short of I) investigating the impact of multi-granular interactions across recent snapshots, and II) harnessing the expressive semantics of significant links accorded with queries throughout the entire history, particularly events exerting a profound impact on the future. These inadequacies restrict representation ability to reflect historical dependencies and future trends thoroughly. To overcome these drawbacks, we propose an innovative TKG reasoning approach towards Historically Relevant Events Structuring (HisRES). Concretely, HisRES comprises two distinctive modules excelling in structuring historically relevant events within TKGs, including a multi-granularity evolutionary encoder that captures structural and temporal dependencies of the most recent snapshots, and a global relevance encoder that concentrates on crucial correlations among events relevant to queries from the entire history. Furthermore, HisRES incorporates a self-gating mechanism for adaptively merging multi-granularity recent and historically relevant structuring representations. Extensive experiments on four event-based benchmarks demonstrate the state-of-the-art performance of HisRES and indicate the superiority and effectiveness of structuring historical relevance for TKG reasoning.
Index Terms:
@@ Abstract
I Introduction
@@
Temporal Knowledge Graphs (TKGs) [1] have revolutionized the limitations of static KGs by considering temporal information. Specifically, TKGs advance beyond traditional triple facts, enhancing them into (subject, relation, object, timestamp) quadruples, thereby bridging the divide between static and dynamic characteristics. This enrichment facilitates capturing the dynamic evolution of real-world events and holds significant practicality. It expands the downstream applications to include event prediction [2], temporal question answering [3], social network analysis [4], anomaly detection [5], etc.
As established in prior research [6, 7], TKGs manifest in the form of sequential snapshots or subgraphs partitioned along a timeline, where each snapshot encapsulates concurrent events occurring at specific timestamps, thereby providing an intuitive representation of event evolutionary patterns. However, TKGs persistently suffer from incompleteness and encounter unique challenges attributable to their temporal characteristics. The critical task of addressing these limitations can be categorized into interpolation and extrapolation, involving reasoning missing facts (events) at past and future timestamps according to historical knowledge, respectively.
From the perspective of history utilization, extrapolation methods can be classified into two principal categories: a) tallying statistics of historical events and leveraging repetitive patterns [8, 9]; b) capturing concurrent and evolutionary correlations from the most recent snapshots [10, 6, 11, 12, 13, 14]. In detail, the a) category generates sparse matrices from repetitive statistics to mask predictions. The b) category adopts Graph Neural Networks (GNNs) to model structural dependencies among concurrent facts, and Recurrent Neural Networks (RNNs) to capture the sequential evolution of snapshots. Based on the b) category, few innovative models [7, 15, 16] are proposed to leverage global historical information. Among them, TiRGN [7] employs a global historical encoder to constrain the prediction scope. HGLS [15] establishes connections between identical entities across different timestamps to capture multi-hop correlations among globally historical events. LogCL [16] constructs a graph consisting of historical facts relevant to queries to obtain rich semantics.
Despite these advancements, previous studies still fall short of comprehensively exploring historically relevant events in TKGs. This limitation stems from two specific challenges: I) Recent Event Impact: investigating the impact of multi-granular interactions across recent snapshots, and II) Distant Historical Influence: harnessing expressive semantics of significant links accorded with queries throughout the entire history, particularly events exerting profound future impact.
I) Recent Event Impact: the first noteworthy challenge arises from inadequate explicit and flexible assessment of multi-granular correlations between facts across recent adjacent snapshots. Recent facts typically exert stronger influence on future events and can initiate chains of events spanning multiple temporal intervals. As sequentially occurring events at adjacent timestamps often share causal relations, the aforementioned deficiency results in the loss of crucial inter-event dependencies. Although RPC [13] and LogCL [16] incorporate inter-snapshot information and assign variable weights to differentiate contributions of recent snapshots, they remain limited to the individual snapshot level, potentially overlooking fact-level propagation patterns that provide critical signals for effective TKG reasoning.
II) Distant Historical Influence: the second significant challenge concerns temporally distant historical facts that can substantially influence future events despite their remoteness. Existing works frequently overlook crucial information intrinsic in globally relevant event structures, particularly long-term semantic patterns and periodic interactions, resulting in inadequate representations. While HGLS [15] attempts to address this by connecting identical entities across snapshots through additional edges, it incorporates redundant information from distant timestamps and requires excessive computational complexity. LogCL [16] integrates query-relevant historical facts but fails to explicitly prioritize key historical links during aggregation, thereby limiting its ability to differentiate the relative importance of historical facts from various timestamps.
Figure 1: Example of extrapolation over TKGs. Left: original historical snapshots containing timestamped facts. Right: derived graphs structured through multi-granularity correlation (red elements) and global relevance (blue elements) perspectives.
To tackle these aforementioned challenges, this paper classifies them into two specific interactions within TKGs: multi-granularity correlation and global relevance. Figure 1 exemplifies these concepts using an instance from ICEWS [17]. Each entity is assigned a unique numeric identifier (e.g., $e_{1}$ denotes North_America). For clarity, the entities on the right side of Figure 1 are represented as simplified nodes.
  • Multi-granularity correlation (denoted by red dotted frame) captures inter-snapshot correlations by structuring facts across varying span of recent timestamps, rather than complex evolutionary patterns. The nodes and edges (red arrow) represent the aggregation of all facts from two adjacent snapshots. Such structure enables the capture of lasting effects through two-hop paths across consecutive timestamps (e.g., $e_{2}\rightarrow r_{2}\rightarrow e_{1}\rightarrow r_{1}\rightarrow e_{3}$, where $(e_{2},r_{2},e_{1})$ occurs at $t-2$, and $(e_{1},r_{1},e_{3})$ at $t-1$). This addresses the challenge of I) Distant Historical Influence by highlighting the interplay between facts in the most recent snapshots.
  • Global relevance (denoted by blue solid frame) encompasses significant facts related to the query ($e_{1}$, $r_{1}$) by structuring links (blue arrow) throughout the original historical snapshots. This addresses the challenge of II) Distant Historical Influence, emphasizing key events that hold profound implications for future from distant timestamps among numerous historical events. From a global relevance perspective, the occurrence of $(e_{1},r_{1},e_{0})$ at $t=0$ may influence $(e_{1},r_{1},e_{3})$ at $t-1$ through two-hop path connections.
@@
Focusing on these two interactions, this paper proposes a novel method that specializes in TKG reasoning with Historically Relevant Event Structuring (HisRES). Specifically, HisRES adopts an encoder-decoder architecture, focusing on two critical aspects of TKG reasoning: capturing the structural and temporal patterns within recent snapshots and understanding the impact of relevant events throughout the entire history. In terms of recent history, HisRES employs a multi-granularity evolutionary encoder (§III-B), which models concurrent interactions and evolution within each snapshot and further integrates adjacent snapshots to learn temporal dependencies using multi-hop links (addressing the first challenge). To explicitly capture structural correlations across the entire history (addressing the second challenge), HisRES introduces a global relevance encoder, structuring historical facts directly relevant to the query set. We propose an innovative convolution-based graph attention network (ConvGAT) to efficiently emphasize important links of historical relevance for future prediction. Furthermore, HisRES incorporates a self-gating mechanism (§III-C) to adaptively fuse entity representations from different granularities or encoders, achieving a comprehensive understanding by considering various historical information (i.e., global and local, intra-snapshot and inter-snapshot) of TKGs, leading to accurate predictions. Extensive experiments on four benchmarks demonstrate the superiority and effectiveness of HisRES in TKG reasoning, particularly in structuring historical relevance.
@@
  • We propose a novel TKG reasoning method termed HisRES, which sufficiently discovers structural and temporal dependencies through historically relevant event structuring.
@@
  • Through the adaptive fusion of multi-granularity evolutionary and global relevance encoding, HisRES can effectively integrate representations throughout historical interactions.
@@
  • Extensive experiments on four event-based datasets demonstrate that HisRES outperforms state-of-the-art TKG reasoning baselines, showcasing its effectiveness and superiority.
II Related Work
@@
KG reasoning [18] has been extensively studied in recent years, making significant contributions to downstream applications. Traditional reasoning deals mainly with static KGs such as the widely used FB15k-237 [19] and WN18RR [20]. In contrast, TKG reasoning is tailored for time-sensitive data such as ICEWS [17], adept at capturing evolutionary patterns, and discovering future facts within event-based datasets characterized by temporal features. Our related work encompasses both static and temporal knowledge graph reasoning.
II-A Static Knowledge Graph Reasoning
@@
Early research on static KGs focuses on modeling entities and relations in various representation spaces. TransE [21] introduces a simple translation assumption $s+r\approx o$, while DistMult [22] learns relational semantics through embeddings from a bilinear objective. In the complex vector space, ComplEx [23] and RotatE [24] employ Hermitian dot products and relation rotations, respectively. With the advent of convolutional neural networks, approaches such as ConvE [20] and ConvTransE [25] utilize convolutional operations to fuse entity-relation embeddings. Recent advancements aim at learning more expressive structures with Graph Convolutional Networks (GCN) [26]. R-GCN [27] extends GCNs with relation-specific block-diagonal matrices. CompGCN [28] integrates multi-relational GCN methods with various composition operations. To differentiate node contributions, GAT [29] employs attention mechanisms, while KBGAT [30] uses relation-specific attention to capture neighboring semantics. However, these static approaches exhibit critical limitations: they neither effectively incorporate explicit temporal information nor adequately explore complex periodicity and historical influences. These deficiencies significantly constrain their ability to handle dynamic facts and learn evolutionary patterns.
II-B Temporal Knowledge Graph Reasoning
@@
TKG reasoning can be broadly categorized into two types based on the occurrence moment of predicted events: interpolation and extrapolation. The former deals with known timestamps, while the latter addresses future timestamps. Additionally, a few methods [31, 32, 33] aim to predict time through inductive reasoning based on logical rules. Notably, this paper focuses on the extrapolation.
Assuming we know all facts before $t_{T}$ (including $t_{T}$), the extrapolation task aims to predict future facts at timestamps $t>t_{T}$ based on historical TKG sequences. Various approaches have been explored in recent years to address it. xERTE [34] designs a subgraph sampling module based on temporal information and incorporates time within relation embeddings. GHNN [35] proposes graph hawkes neural network to capture the evolution of graph sequences. TANGO [36] extends neural ordinary differential equations to model dynamic knowledge, enabling continuous-time representation. TITer [37] and CluSTER [38] employ reinforcement learning to generate query-specific paths composed of historical entities. TLogic [39] incorporates symbolic representations and learns temporal logical rules to enhance reasoning.
@@
Besides, numerous studies can be categorized into two main groups: those that focus on repeating statistical historical facts, and those that model the interactions among historical events. In the first category, CyGNet [8] employs a multi-hot vector as a historical vocabulary to record past facts, subsequently using this information to mask predictions. Building upon this, CENET [9] and PLEASING [40] introduce contrastive learning and learnable classifiers to differentiate historical and potential events. In the latter category, RE-NET [10] and RE-GCN [6] employ GCNs to model concurrent structures and utilize recurrent units to learn entity or relation evolution patterns. On this basis, CEN [11] considers historical length-diversity and time-variability. TiRGN [7] adopts a global history encoder like CyGNet to identify entities related to historical or unseen facts and a time-guided decoder. RETIA [12] and RPC [13] introduce the concept of line graphs to learn the interaction of relations. LMS [41] learns entity and time interactions separately from a multi-graph perspective. L2TKG [14] addresses the missing associations and creates links between unconnected entities to obtain more comprehensive representations. Some novel insights attempt to capture global structural correlations through additional graphs. HGLS [15] explicitly captures long-term dependencies by connecting the same entity at each timestamp, while LogCL [16] proposes an entity-aware attention mechanism to flexibly adopt facts relevant to queries and utilizes contrastive learning to improve robustness.
Nevertheless, existing methods still face limitations in capturing two key interactions: I) Distant Historical Influence and II) Distant Historical Influence. Regarding the former, although RPC and LogCL propose mechanisms to calculate the correspondence of recent timestamps and distinguish the importance of different recent snapshots respectively, they fail to explicitly model the impact propagation of individual facts from the most recent snapshots—a capability crucial for predicting future repetitive or unseen facts. Concerning the latter, TiRGN employs a fixed history mask matrix to focus on historical repetitive facts, but this constrains the model’s predictive capability as it fails to learn information from distant historical data. While HGLS establishes edges between identical entities at different timestamps to capture historical associations through multi-hop relations, it significantly multiplies the graph’s node count and necessitates a multi-layer to capture multi-hop information, leading to increased computational overhead. Additionally, LogCL suffers from inefficient encoding when aggregating global information, making it difficult to differentiate the contributions of historical facts from different temporal periods.
III Method